SwaggyMark:

A plane traveled 1512 miles each way to Athens and back. The trip there was with the wind. It took 9 hours. The trip back was into the wind. The trip back took 12 hours. Find the speed of the plane in still air and the speed of the wind.

Rx1nyDxyz:

Ok do you have any idea of where to start here?

surjithayer:

let the speed of plane in still air=x mph speed of wind =y mph speed of plane with wind=d/t=1512/9=168 mph speed of plane into wind=1512/12=126 mph x+y=168 x-y=126 adding 2x=294 x=147 mph 147+y=168 y=168-147=21 mph speed of plane in still air=147 mph speed of wind=21 mph